Ivy went for a trip and spent 3/5 of her answer on gas for her car. She spent 1/2 of the remaining money for lunch. She spent Php250.00 for her lunch. How much money does she had before she left?
Answer by Theo(13342)

let x equal the money she had before she left.

she spent 3/5 * x on gas.

this means she had 2/5 * x left

she spent 1/2 of that on lunch.

this means she spent 1/2 * 2/5 * x = 2/10 * x = 1/5 * x on lunch

lunch cost 250, therefore 1/5 * x = 250.

your equation to solve is 1/5 * x = 250

multiply both sides of that equation by 5 to get:

1/5 * 5 * x = 5 * 250

simplify to get x = 1250

that's what she had before she left.

she had 1250 before she left.

she spent 3/5 of that on gas.

she spent 3/5 * 1250 = 750 on gas.

this means she had 1250 - 750 = 500 left.

she spent half of that on lunch.

this means she spent 250 on lunch.

the numbers check out.

your solution is that she had 1250 before she left on her trip.